# Invertible Calabi-Yau Orbifolds over Finite Fields

### v1

@@ 4. An Arithmetic Conjecture
$x_{2}^{1}x_{4}^{10}y_{1}^{1}y_{3}^{5}\theta_{2}\theta_{4}$ $-p\,\Gamma_{p}\left(\frac{1}{3}\right)\Gamma_{p}\left(\frac{2}{3}\right)$
$x_{2}^{2}x_{4}^{5}y_{1}^{1}y_{3}^{5}\theta_{2}\theta_{4}$ $-p\,\Gamma_{p}\left(\frac{2}{3}\right)\Gamma_{p}\left(\frac{1}{3}\right)$
$x_{1}^{1}x_{3}^{5}y_{2}^{1}y_{4}^{10}\theta_{1}\theta_{3}$ $-p\,\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{1}{2}\right)$
$x_{1}^{1}x_{3}^{5}y_{2}^{2}y_{4}^{5}\theta_{1}\theta_{3}$ $-p\,\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{1}{2}\right)$
@@
$\left(0,2\right)$ $x_{1}^{1}x_{2}^{2}x_{3}^{9}x_{4}^{14}\theta_{1}\theta_{2}\theta_{3}\theta_{4}$ $-p^{2}\,\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{2}{3}\right)\Gamma_{p}\left(\frac{9}{10}\right)\Gamma_{p}\left(\frac{14}{15}\right)$
@@ Acknowledgements
It is a pleasure to acknowledge insightful exchanges with Chuck Doran, Doug Haessig, Tristan Hübsch, Tyler Kelly, Bong Lian, Ursula Whitcher, and Noriko Yui over the long gestation time of this work. M.A. was supported in part by VCU Quest Award “Quantum Fields and Knots: An Integrative Approach.” Some of the work was carried out while A.P. visited VCU in connection with the Richmond Geometry Meeting funded by NSF grant DMS-2349810, and while A.P. held a Coleman Postdoctoral Fellowship at Queens University, supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada through the Discovery Grant of Noriko Yui.

### v2

@@ 4. An Arithmetic Conjecture
$x_{2}^{1}x_{4}^{10}y_{1}^{1}y_{3}^{5}\theta_{2}\theta_{4}$ $p\,\Gamma_{p}\left(\frac{1}{3}\right)\Gamma_{p}\left(\frac{2}{3}\right)$
$x_{2}^{2}x_{4}^{5}y_{1}^{1}y_{3}^{5}\theta_{2}\theta_{4}$ $p\,\Gamma_{p}\left(\frac{2}{3}\right)\Gamma_{p}\left(\frac{1}{3}\right)$
$x_{1}^{1}x_{3}^{5}y_{2}^{1}y_{4}^{10}\theta_{1}\theta_{3}$ $p\,\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{1}{2}\right)$
$x_{1}^{1}x_{3}^{5}y_{2}^{2}y_{4}^{5}\theta_{1}\theta_{3}$ $p\,\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{1}{2}\right)$
@@
$\left(0,2\right)$ $x_{1}^{1}x_{2}^{2}x_{3}^{9}x_{4}^{14}\theta_{1}\theta_{2}\theta_{3}\theta_{4}$ $p^{2}\,\Gamma_{p}\left(\frac{1}{2}\right)\Gamma_{p}\left(\frac{2}{3}\right)\Gamma_{p}\left(\frac{9}{10}\right)\Gamma_{p}\left(\frac{14}{15}\right)$
@@ Acknowledgements
It is a pleasure to acknowledge insightful exchanges with Chuck Doran, Doug Haessig, Tristan Hübsch, Tyler Kelly, Bong Lian, Daqing Wan, Ursula Whitcher, and Noriko Yui over the long gestation time of this work. M.A. was supported in part by VCU Quest Award “Quantum Fields and Knots: An Integrative Approach.” Some of the work was carried out while A.P. visited VCU in connection with the Richmond Geometry Meeting funded by NSF grant DMS-2349810, and while A.P. held a Coleman Postdoctoral Fellowship at Queens University, supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada through the Discovery Grant of Noriko Yui.